OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, –2) and has a slope of –3?

OpenStudy (anonymous):

y - b = m(x - a) y=y b=-2 m=-3 x=x a=4

OpenStudy (anonymous):

those arent the choices x + 3y = 10 x + 3y = –10 3x + y = 10 3x + y = –10

OpenStudy (anonymous):

I'm not giving you the answer, you can work it out yourself with what I've given you...

OpenStudy (anonymous):

y=mx + c (4,-2) and m=-3 -2=-3(4) + c c=..... now re-write your whole equation with the gradient and y-intercept.

OpenStudy (anonymous):

i dont get it i dont get any answer that is on their are you sure your doing it the right way?

OpenStudy (anonymous):

y + 2 = -3(x-4) y + 2 = -3x + 12 Rearrange and you'll get your answer

OpenStudy (anonymous):

@ScottB05 im not asking for the answers but can you show me the steps cus im not understanding any of this

OpenStudy (anonymous):

yeh, doing @ScottB05 or my way, you should get the same answer. I think you just have a problem with rearranging the equation.

OpenStudy (anonymous):

Well okey, the standard formula for a linear equasion is; y = m x + c Y and X are your co-ordinates, M is your gradient or 'slope' and c is your constant. Now, lets do it Jay's way as it's probebly easier to understand. Sup in x and y (4, -2) -2 = 4m + c Now, lets sub in m, -3 -2 = -12 + c So c = -2 + 12...................c = 10 Now go abck up to the top... Sub in c and m this time to the linear equasion y=mx+c y = -3x + 10 Rearrange this; y + 3x = 10

OpenStudy (anonymous):

y= 3x - 10 ?

OpenStudy (anonymous):

It's -3

OpenStudy (anonymous):

And it's + 10

OpenStudy (anonymous):

@ScottB05 these are the choices x + 3y = 10 x + 3y = –10 3x + y = 10 3x + y = –10 i think its D

OpenStudy (anonymous):

y + 3x = 10 is what I got and that's there at C... But you can work through my working again and see if I made a mistake, but until then I'm pretty confident about it being C

OpenStudy (anonymous):

Okay im going to review what you put and thank you for your time and patience :)

OpenStudy (anonymous):

No problem :) Good luck ;D